The following is the sales budget for Profit, Inc., for the first quarter of 2018:

January February March
  Sales budget $ 216,000 $ 236,000 $ 259,000

Credit sales are collected as follows:

65 percent in the month of the sale
30 percent in the month after the sale
5 percent in the second month after the sale

The accounts receivable balance at the end of the previous quarter was $100,000 ($70,000 of which was uncollected December sales).

a. Calculate the sales for November. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
b. Calculate the sales for December. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
c. Calculate the cash collections from sales for each month from January through March. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1
a.
Uncollected November sales at the end of previous quarter = Accounts receivables balance at the end of previous quarter - Uncollected December sales = 100000 - 70000 30000
Sales for November = Uncollected November sales at the end of previous quarter / % Uncollected = 30000 / 5% 600000.00
b.
Sales for December = Uncollected December sales at the end of previous quarter / % Uncollected = 70000 / 35% 200000.00
c.
Cash collections from sales : January February March

65% of current month sales

140400

[ 216000*65% ]

153400

[ 236000*65% ]

168350

[ 259000*65% ]

30% of previous months sales

60000

[ 200000*30% ]

64800

[ 216000*30% ]

70800

[ 236000*30% ]

5% of sales made 2 month ago

30000

[ 600000*5% ]

10000

[ 200000*5% ]

10800

[ 216000*5% ]

Cash collections 230400.00 228200.00 249950.00